### v1

# Instructions for *ACL Proceedings

Abstract
Counter-speech (CS) is a key strategy for mitigating online Hate Speech (HS), yet defining the criteria to assess its effectiveness remains an open challenge. We propose a novel computational framework for CS effectiveness classification, grounded in social science concepts. Our framework defines six core dimensions – Clarity, Evidence, Emotional Appeal, Rebuttal, Audience Adaptation, and Fairness – which we use to annotate 4,214 CS instances from two benchmark datasets, resulting in a novel linguistic resource released to the community. In addition, we propose two classification strategies, multi-task and dependency-based, achieving strong results (0.94 and 0.96 average F1 respectively on both expert- and user-written CS), outperforming standard baselines, and revealing strong interdependence among dimensions.
Effectiveness of Counter-Speech against Abusive Content: A Multidimensional Annotation and Classification Study
Greta Damo1, Elena Cabrio1 Serena Villata1 1Université Côte d’Azur, CNRS, Inria, I3S, France Correspondence: greta.damo@univ-cotedazur.fr
Content warning: this paper contains unobfuscated examples some readers may find offensive.
1 Introduction
The rise of social media has intensified concerns over online Hate Speech (HS). Counter-speech (CS), non-aggressive responses that challenge hate using facts or alternative views, has emerged as a promising mitigation strategy Schieb and Preuss (2016); Benesch (2014). CS has been studied through various lenses: some see it as direct responses to hate Mathew et al. (2019); Ashida and Komachi (2022), while others argue it should condemn hate and offer victims support and alternative views Vidgen et al. (2021); Hangartner et al. (2021); Vidgen et al. (2020); He et al. (2021). CS is also valued for its potential to shift beliefs and reduce future hate Qian et al. (2019a); Rieger et al. (2018). CS can be written by experts, online users, or chatbots (based on Large Language Models – LLMs). Expert-written CS typically presents structured reasoning and appeals to shared human values, whereas user and LLM responses tend to be more generic and emotional Mun et al. (2023); Tekiroğlu et al. (2022, 2020).
This study evaluates expert- and user-written CS effectiveness using two benchmark datasets Albanyan and Blanco (2022); Chung et al. (2019). Figure 1 shows examples of expert and user CS.
Figure 1: Examples of expert-written and user-written counter-speech.
Current evaluations on CS effectiveness rely on surface-level metrics – coherence, relevance, or non-offensiveness – assuming well-formed CS is inherently effective Chung et al. (2021); Zhu and Bhat (2021); Baheti et al. (2021); Ashida and Komachi (2022); Mun et al. (2023). But concentrating on CS surface form only is not enough to achieve shifting attitudes or reducing harm.
To address this gap, our contribution is twofold: (1) a new multi-dimensional framework for automatic CS effectiveness classification across six dimensions – Clarity, Evidence, Emotional Appeal, Rebuttal, Audience Adaptation, and Fairness – from linguistics and argumentation theories, and two classification strategies, multi-task and dependency-based modeling, outperforming standard baselines, achieving average F1 scores of 0.94 and 0.96 respectively, when combining user-written and expert-written CS data. (2) a novel human annotated resource combining two expert- Chung et al. (2019) and user-written Albanyan and Blanco (2022) CS datasets, enriched with the annotation layer with the six dimensions [^1] Datasets, guidelines, and code are available at this link..
2 Related Work
CS, introduced as a mitigation strategy by Susan et al. (2016), has not yet been extensively explored from a detection standpoint. Among the studies, Poudhar et al. (2024) find that a one-shot prompted LLM achieves promising accuracy in classifying manually labeled CS strategies. Mathew et al. (2018) analyze HS/CS tweet pairs, using a boosting algorithm with TF-IDF and lexical features, which achieves 0.77 F1. Mathew et al. (2019) also analyze CS and neutral comments on YouTube, finding that TF-IDF vectors combined with logistic regression achieve 0.73 F1. Ziems et al. (2020) analyze tweets during the COVID-19 crisis, achieving 0.49 F1 for CS. Garland et al. (2020) apply an ensemble learning algorithm to classify HS and CS tweets, achieving F1 scores between 0.76 and 0.97. A key gap in the literature is the lack of standardized evaluation methods for assessing CS. Several studies propose evaluation criteria, but no unified framework emerged. Zheng et al. (2023) assess CS effectiveness through human judgments, focusing on comfort for targets and bystander empathy. Chung et al. (2021) and Zhu and Bhat (2021) similarly rely on human ratings based on qualities such as informativeness, coherence, and stance. Ashida and Komachi (2022) evaluate CS suitability using offensiveness and stance, while Jones et al. (2024) propose an LLM-based evaluation framework, grounded in NGO guidelines, which shows strong alignment with human judgments. Yu et al. (2024) introduce a conversation-level incivility metric, arguing that CS should be judged by its downstream impact on discourse. We advance beyond previous studies by introducing a human-annotated resource that captures fine-grained dimensions of structural and rhetorical message quality, and by developing classification methods that leverage this framework.
3 Effectiveness Metrics
We propose a framework to evaluate CS effectiveness across six dimensions: Clarity, Evidence, Emotional Appeal, Rebuttal, Audience Adaptation, and Fairness. Clarity ensures structural coherence and logical flow, allowing to follow texts effectively Persing and Ng (2013); Wachsmuth et al. (2016, 2017); Stapleton and Wu (2015). Evidence strengthens persuasiveness through relevant support such as statistics or expert testimony Wachsmuth et al. (2017); Rinott et al. (2015); Rahimi et al. (2014); Reinard (1988). Emotional Appeal, a rhetorical strategy, shows that empathy and emotion can enhance impact and share Wachsmuth et al. (2017); Aristotle (2007); Govier (2010); Hangartner et al. (2021); Chen et al. (2022); Marone (2015). Rebuttal reflects the ability to address and counter opposing views, improving credibility and persuasion Wachsmuth et al. (2017); Aristotle (2007); Toulmin (1958); Damer (2009); Granger et al. (2009); O’Keefe (1999); Onoda et al. (2015). Audience Adaptation involves tailoring language to the audience’s style and knowledge, improving relatability Pickering and Garrod (2004); Danescu-Niculescu-Mizil et al. (2011); Wang et al. (2014). Fairness promotes respectful tone and avoids dehumanization, supporting ethical discourse Cialdini (1993); Schreier et al. (1995).
Categorical Dimensions (1–3 Likert Scale)
Clarity 3: Clear, logically structured, directly addresses the HS topic.
2: Mostly clear or slightly generic.
1: Vague or incoherent.
Evidence 3: Multiple pieces of information as supporting evidence.
2: One supporting information.
1: No information as evidence.
Rebuttal 3: Multiple rebuttals targeting specific parts of the HS.
2: One rebuttal.
1: No rebuttal.
Fairness 3: Respectful, no swearing, no attacks against the hater.
2: Mostly polite, no attacks, mild offensive.
1: Aggressive, including swearing and personal attacks.
Binary Dimensions (0/1)
Emotional Appeal 1: Language that evokes emotions (either positive or negative).
0: Neutral tone and language.
Audience Adaptation 1: Matches HS tone/complexity.
0: Mismatched HS tone or complexity.
Table 1: Annotation guidelines for CS effectiveness.
Emotional Appeal and Audience Adaptation are binary (0–1) variables, where 1 indicates presence and 0 absence. Similar to prior work simplifying subjective dimensions into binary formats Habernal and Gurevych (2016); Wachsmuth et al. (2017); Oraby et al. (2017); Koszowy et al. (2024), we adopt this approach to reduce ambiguity. Clarity, Evidence, Rebuttal, and Fairness are categorical, using a 1–3 Likert scale, with 3 as the best score. These dimensions rely on more objective indicators (for example, the number of evidence pieces) that capture weak, moderate, and strong instances. Table 1 reports the definitions for the annotation of the effectiveness dimensions.
4 Datasets Description
We extend the annotation on two existing datasets: CONAN Chung et al. (2019) and the Twitter Dataset Albanyan and Blanco (2022). We select CONAN as it is the first expert-curated dataset of HS/CS pairs, widely used as a benchmark. It is a multilingual (English, French, Italian) dataset centered on Islamophobia, containing 4,078 expert-annotated HS/CS pairs. Through translation and paraphrasing, it is expanded to 14,988 pairs. For our experiments, we retain only the English, non-augmented instances, resulting in 3,847 pairs. The Twitter Dataset [^2] Only tweet IDs are publicly available due to privacy policies. We obtained the full data upon request from the authors. contains 5,652 hateful tweets and replies from social media (Twitter/X), capturing the brevity and style typical of online discourse Bonaldi et al. (2024). We focus on the subset labeled as CS, where a clear target is identifiable, yielding 367 HS/CS pairs. Together, these datasets form a combined benchmark of 4,214 HS/CS pairs.
Annotation & IAA. We define annotation guidelines to label CS effectiveness dimensions. To consolidate these guidelines, a pilot study was conducted on 50 HS/CS pairs from each dataset, involving three annotators  [^3] Two female, one other; age group: 21–30; education level: PhD students. with background in computer science. Through different reconciliation phases, the guidelines were improved iteratively. After an initial round, annotators reviewed disagreements to resolve confusion or misinterpretation. Based on their feedback, the guidelines were updated, and a second and final annotation round was conducted. To measure inter-annotator agreement (IAA), Fleiss’s $\kappa$ was used for binary dimensions, Krippendorff’s $\alpha$ for categorical ones, and Percent Agreement for Audience Adaptation due to near-perfect consensus (as other metrics underestimate high non-random agreement). IAA results improved significantly in the second round, after which the final dataset was labeled by one of the annotators. Table 2 shows IAA scores, indicating strong agreement across all dimensions (see Appendix A for pairwise scores). We also show average effectiveness scores, demonstrating that expert-written CS is more effective (avg. 1.62) than user-written CS. CONAN scores higher in all dimensions except emotional appeal, suggesting online users rely more on emotions than evidential reasoning or facts.
Dimension CONAN Twitter
1st 2nd avg. 1st 2nd avg.
emotional 0.30 0.65 0.29 0.40 0.62 0.45
audience 1.00 1.00 0.99 0.99 0.99 1.00
clarity 0.29 0.75 2.72 0.41 0.82 1.96
evidence 0.27 0.73 1.47 0.34 0.75 1.07
rebuttal 0.21 0.78 1.31 0.42 0.62 1.18
fairness 0.63 0.79 2.95 0.53 0.94 2.35
Total AVG 1.62 1.34
Table 2: IAA from two annotation rounds (Fleiss’ $\kappa$ for binary variables, and Krippendorff’s $\alpha$ for categorical ones) and average effectiveness scores.
5 Experimental Setting
We use BERT-based classifiers to predict CS effectiveness scores across the six dimensions, relying on bert-base-uncased Devlin et al. (2018) trained on the combined dataset of 4,214 HS/CS pairs, and cross-dataset evaluations. We use BERT as the sole model to focus on improving it through our framework. By not comparing across different architectures, we isolate the effect of our multi-task and dependency modeling, ensuring that observed gains result from our contributions. We assess the following configurations (Appendix B for details):
Bert_CS: CS text embeddings are used as input. Binary dimensions are trained using Binary Cross-Entropy (BCE) loss function, and categorical ones with Cross-Entropy (CE). Bert_CS_HS: identical to Bert_CS, but the input is the concatenation of HS and CS texts embeddings for more context. Multi-task_divided trains binary dimensions jointly (BCE for Emotional Appeal, Focal Loss for Audience Adaptation) and categorical ones separately (CE), assuming partial correlation. Multi-task_united combines all dimensions in a single model using the sum of their respective loss functions, assuming stronger interdependence. Dependency_matrix_3e introduces a learnable matrix to capture pairwise relationships between dimensions, and is trained for 3 epochs, while Dependency_matrix_6e for 6 epochs to assess the effect of extended training.
6 Results
Model Emo. Aud. Clarity Evid. Rebut. Fair. AVG
Train: Combined Dataset
bert_cs 0.67±0.09 0.99±0.00 0.94±0.01 0.96±0.01 0.96±0.01 0.95±0.01 0.91±0.02
bert_cs_hs 0.62±0.12 0.99±0.00 0.94±0.02 0.96±0.01 0.95±0.01 0.94±0.01 0.90±0.02
multi-task_d 0.92±0.02 0.99±0.00 0.92±0.01 0.93±0.01 0.94±0.03 0.92±0.03 0.94±0.01
multi-task_u 0.84±0.04 0.99±0.00 0.93±0.02 0.92±0.02 0.93±0.03 0.93±0.02 0.92±0.01
dep._m._3e 0.80±0.10 0.99±0.00 0.92±0.01 0.90±0.05 0.91±0.03 0.93±0.03 0.91±0.02
dep._m._6e 0.92±0.04 0.99±0.00 0.95±0.01 0.97±0.02 0.96±0.01 0.95±0.02 0.96*±0.01
Train: CONAN $\rightarrow$ Test: Twitter
bert_cs 0.30±0.01 1.00±0.00 0.41±0.04 0.86±0.09 0.77±0.02 0.54±0.21 0.65±0.05
bert_cs_hs 0.30±0.04 1.00±0.00 0.33±0.03 0.84±0.07 0.73±0.05 0.41±0.07 0.60±0.02
multi-task_d 0.49±0.07 1.00±0.00 0.42±0.02 0.91±0.04 0.80±0.02 0.32±0.03 0.65±0.02
multi-task_u 0.52±0.06 1.00±0.00 0.42±0.03 0.93±0.04 0.77±0.02 0.35±0.05 0.66±0.01
dep._m._3e 0.54±0.02 1.00±0.00 0.41±0.02 0.92±0.02 0.77±0.02 0.34±0.05 0.66±0.01
dep._m._6e 0.50±0.05 1.00±0.00 0.42±0.01 0.92±0.03 0.77±0.02 0.38±0.01 0.66*±0.01
Train: Twitter $\rightarrow$ Test: CONAN
bert_cs 0.14±0.00 0.99±0.00 0.74±0.05 0.45±0.06 0.55±0.11 0.92±0.01 0.63±0.03
bert_cs_hs 0.14±0.01 0.99±0.00 0.76±0.06 0.43±0.02 0.55±0.14 0.92±0.02 0.63±0.03
multi-task_d 0.55±0.17 0.99±0.00 0.51±0.24 0.44±0.01 0.60±0.05 0.89±0.13 0.66*±0.04
multi-task_u 0.47±0.23 0.99±0.00 0.38±0.27 0.43±0.02 0.56±0.07 0.72±0.39 0.59±0.09
dep._m._3e 0.40±0.24 0.99±0.00 0.48±0.22 0.43±0.03 0.53±0.09 0.70±0.13 0.59±0.08
dep._m._6e 0.40±0.21 0.99±0.00 0.75±0.04 0.53±0.07 0.47±0.21 0.78±0.14 0.65±0.07
Table 3: F1 scores (± std) for all models with CS embeddings. Statistically significant results are marked with *.
Table 3 shows average F1 scores across five runs (seeds 42, 0, 1, 2, 3) for all model configurations across our CS effectiveness dimensions. We compare BERT baselines, multi-task learning models, and dependency-based architectures (with CS embeddings). Dependency_matrix_6e consistently achieves the highest average F1 score (0.96), outperforming all the others. This performance gain is likely due to the use of a learnable dependency matrix that captures the interrelations between dimensions, allowing the model to exploit cross-dimensional dependencies, unlike models that treat the six dimensions independently. Statistical tests confirm the results differ significantly, especially between the best model, Dependency_m._6e, and all others (Appendix C). Figure 2 visualizes the dependency matrix learned by Dependency_m._6e. Rows represent influenced tasks, columns represent influencing tasks, with higher values indicating stronger influence. Emotional Appeal, for instance, is influenced by Fairness (0.63) and Rebuttal (0.65), while Audience Adaptation is shaped by Clarity (0.98), Rebuttal (0.96), and Fairness (0.84). These inter-dependencies help the model to improve its classification performance across dimensions. Both multi-task models, Multi-task_d. (0.94) and Multi-task_u. (0.92), outperform the BERT baselines, suggesting that the multi-task framework is more effective, with only a slight performance difference between united and divided. Incorporating HS data in Bert_CS_HS (0.90) does not significantly improve performance over Bert_CS (0.91). Emotional Appeal remains the hardest dimension to classify, showing the most variation across models. In this dimension, dependency-based and multi-task models perform best (up to 0.92 for Dependency_m._6e and Multi-task_d.), while BERT baselines lag behind (0.67 and 0.62), suggesting that richer architectures better capture emotional cues. Audience Adaptation scores above 0.99 in all configurations (Appendix E).
Figure 2: Learned Dependency Matrix obtained from dep._m._6e on the combined dataset with seed 42.
Cross-Domain Generalization. In the CONAN $\rightarrow$ Twitter setting, all models show a marked performance drop, with Dependency_m._6e performing best (0.66) alongside Dependency_m._3e and Multi-task_u. Emotional Appeal declines the most, suggesting a domain shift in how emotional content is conveyed. Clarity, Fairness, and Rebuttal scores are also low, reflecting the less structured nature of user-written CS. In the reverse setting, Twitter $\rightarrow$ CONAN, results are similar. Multi-task_d. achieves the highest avg. F1 score (0.66), followed closely by Dependency_m._6e (0.65). The remaining models obtain lower average F1 scores, indicating less robust generalization in this transfer setup. Models trained on expert-written CS seem more adaptable, but Emotional Appeal remains challenging across domains, ranging from 0.14 to 0.55. In contrast, Audience Adaptation remains consistently high. In-Domain Evaluation. Training and testing on the same dataset (Appendix D) confirm that dependency-based models consistently perform best, further supporting the benefit of modeling interrelations between dimensions in classification.
7 Conclusion
As a first contribution of this paper, we release a novel linguistic resource that merges the CONAN and Twitter datasets for CS, manually annotated with six dimensions of CS effectiveness. As a second contribution, we present a framework for automatically classifying CS effectiveness across these dimensions. Results show that the dependency-based model outperforms BERT baselines with 0.96 F1 on both expert- and user-written CS. Multi-task models also surpass BERT with 0.94 F1, confirming that effectiveness dimensions are interdependent and benefit from explicit dependency modeling. Our architecture-agnostic framework improves BERT and generalizes to other transformers. A deep linguistic analysis highlights which dimensions need improvement for CS evaluation.
8 Limitation
Our framework for assessing CS effectiveness dimensions focuses on key aspects such as structural coherence, linguistic strategies, content strength, and rhetorical impact. Given the vast body of literature on counter-speech in social sciences, we focused on the most relevant studies, but acknowledge that further conceptualizations exist.
Another limitation concerns validating the effectiveness of CS with human users. Future work should assess whether CS aligned with our dimensions effectively influences discourse, fosters empathy toward targets, induces belief change in users, and reduces hateful content from the original authors.
Regarding the data, we acknowledge that our dataset is imbalanced, with a higher number of counter-speech examples generated by experts compared to those from online users, a common trait in counter-speech corpora due to the scarcity of high-quality, user-generated responses Susan et al. (2016); Mathew et al. (2019); Chung et al. (2019); Qian et al. (2019b). Additionally, some dimensions, particularly Audience Adaptation, exhibit class imbalance, which may affect model performance. To improve robustness, future studies should leverage larger and more balanced datasets.
Despite these limitations, our study introduces a new evaluation framework for counter-speech effectiveness and provides a model setup that achieves strong classification results. With this work, we aim to support and strengthen the community effort towards the refinement of counter-speech evaluation frameworks and the improvement of automatic approaches for counter-speech effectiveness assessment and generation.
9 Ethical Statement
We are aware that our study involves human exposure to hateful and potentially upsetting content. For this reason, we followed the mitigation strategies proposed by Vidgen et al. (2019) to protect annotators’ mental health. Participants gave their consent to voluntarily participate in the annotation, and were informed that they could withdraw from the study at any time. Moreover, each participant was allocated a small subset of the data (100 examples), and a generous time frame of about three weeks to mitigate prolonged exposure to potentially distressing language. We plan to share the data with the new annotations for research purposes only.
Appendix A Pairwise IAA Scores
Table 4 shows the results of the IAA between each pair of annotators on both CONAN and Twitter datasets, after the first round of annotations. Table 5 shows the pairwise IAA results after the second round. For binary metrics, we use Fleiss’s $\kappa$, while for multi-label metrics we use Krippendorff’s $\alpha$. For the audience adaptation dimension, we use Percent Agreement.
Dimension Ann1 & Ann2 Ann1 & Ann3 Ann2 & Ann3
Tw Co Tw Co Tw Co
emotional_appeal 0.35 0.52 0.51 0.24 0.36 0.18
audience_adaptation 1.00 1.00 0.99 1.00 0.99 1.00
clarity 0.44 0.58 0.56 0.06 0.35 0.30
evidence 0.50 0.70 0.19 0.22 0.21 0.12
rebuttal 0.31 0.32 0.85 0.12 0.28 0.34
fairness 0.56 0.74 0.72 0.57 0.40 0.48
Table 4: First round IAA scores across dimensions for each annotator pair on Twitter (Tw) and CONAN (Co).
Dimension Ann1 & Ann2 Ann1 & Ann3 Ann2 & Ann3
Tw Co Tw Co Tw Co
emotional_appeal 0.88 0.60 0.42 0.85 0.55 0.48
audience_adaptation 1.00 1.00 0.99 1.00 0.99 1.00
clarity 0.88 0.76 0.86 0.79 0.73 0.71
evidence 0.78 0.78 0.85 0.77 0.64 0.62
rebuttal 1 0.72 0.48 0.90 0.48 0.72
fairness 0.92 0.90 0.95 0.78 0.94 0.70
Table 5: Second round IAA scores across dimensions for each annotator pair on Twitter (Tw) and CONAN (Co).
Appendix B Additional Experimental Details
In this section, we describe the model configurations used in our experiments. The goal of these experiments is to classify the effectiveness dimension scores for a given counter-speech response. More specifically, we use bert-base-uncased with 110M parameters. In all configurations, the input length is controlled by setting max_length to 128. The batch_size is 16, the random_state is set to 42, 0, 1, 2, and 3, and the learning_rate is 2e-5. All the inputs are tokenized by using BertTokenizer, and the process includes 3 training epochs (6 for two configurations). For performing the experiments, we concatenate both CONAN and Twitter datasets, and then we split it into train, validation, and test subsets, with a percentage of 70-10-20. We also perform cross-validation, by using a single dataset for training and validation, and the other for testing. The experiments were run on one A100 GPU.
We use the following loss functions:
  • Binary Cross-Entropy Loss (BCE): Used for binary classification tasks, it measures the difference between predicted probabilities and actual labels.
  • Cross-Entropy Loss (CE): Generalized for multi-class classification, CE extends BCE to multiple classes by computing the log loss over softmax-normalized probabilities.
  • Focal Loss: an extension of CE/BCE, focal loss addresses class imbalance by down-weighting easy samples and focusing on hard ones. It introduces a modulating factor to reduce the contribution of well-classified examples.
We also provide more details regarding our model contributions:
  • Multi-task_divided. We take the binary dimensions, and we train bert-base-uncased on them separately, then we combine the two losses (BCE and Focal Loss, for imbalanced datasets) to make the predictions. Then, we take the multi-label dimensions, we train them separately with Cross Entropy Loss, and, at the end, we combine all the losses, and we make the predictions on the test set. This configuration assumes that dimensions, for which the losses are added together, are somehow correlated.
  • Multi-task_united. Here, we take all the dimensions together, both binary and multi-label, and we train them with their respective loss functions (BCE, Focal Loss, and Cross-Entropy loss). Then we aggregate all the losses, and we predict the results for all of them. All predictions are done with the same Multi-task bert model. Again, the configuration assumes that dimensions are somehow correlated.
  • Dependency_learnable. All the dimensions are considered together, by training bert-base-uncased on them. Each dimension has its own loss function (BCE for emotional appeal, Focal Loss for audience adaptation, Cross Entropy Loss for all the multi-label dimensions), which are summed up at the end. In addition, there is a new parameter called dependency-matrix. Its rows and columns contain values between 0 and 1, that are weights accounting for the dependency between pairs of dimensions. Its weights are randomly initialized, and their best value is learned during the training phase. We run the model for 3 and 6 epochs.
We experiment with two types of input embeddings: (i) concatenated Hate Speech (HS) and Counter Speech (CS) embeddings to provide broader context and (ii) CS embeddings alone. The latter consistently yields slightly better performance. Consequently, we adopt CS-only embeddings for all configurations in the main part of the paper (Multi-task_d., Multi-task_u., Dependency_m._3e, and Dependency_m._6e). For completeness, Table 6 also reports the average results over five runs for models using concatenated CS+HS embeddings.
Model Emo. Aud. Clarity Evid. Rebut. Fair. AVG
Train: Combined Dataset
multitask_d. 0.90±0.02 0.99±0.00 0.92±0.01 0.90±0.01 0.91±0.03 0.91±0.03 0.92±0.01
multitask_u. 0.80±0.04 0.99±0.00 0.92±0.02 0.90±0.02 0.91±0.03 0.92±0.02 0.91±0.01
dep._m._3e 0.78±0.10 0.99±0.00 0.91±0.01 0.88±0.05 0.89±0.03 0.93±0.03 0.90±0.02
dep._m._6e 0.90±0.04 0.99±0.00 0.94±0.01 0.96±0.02 0.96±0.01 0.95±0.02 0.95*±0.01
Train: CONAN $\rightarrow$ Test: Twitter
multitask_d. 0.48±0.07 1.00±0.00 0.38±0.02 0.88±0.04 0.80±0.02 0.33±0.03 0.65±0.02
multitask_u. 0.51±0.06 1.00±0.00 0.40±0.03 0.88±0.04 0.79±0.02 0.35±0.05 0.65±0.01
dep._m._3e 0.56±0.02 1.00±0.00 0.39±0.02 0.90±0.02 0.79±0.02 0.34±0.05 0.66*±0.01
dep._m._6e 0.51±0.05 1.00±0.00 0.38±0.01 0.88±0.03 0.79±0.02 0.34±0.01 0.65±0.01
Train: Twitter $\rightarrow$ Test: CONAN
multitask_d. 0.40±0.17 0.99±0.00 0.49±0.24 0.43±0.01 0.60±0.05 0.86±0.13 0.63±0.04
multitask_u. 0.40±0.23 0.99±0.00 0.45±0.27 0.43±0.02 0.54±0.07 0.70±0.39 0.58±0.09
dep._m._3e 0.32±0.24 0.99±0.00 0.39±0.22 0.43±0.03 0.50±0.09 0.73±0.13 0.56±0.08
dep._m._6e 0.35±0.21 0.99±0.00 0.67±0.04 0.48±0.07 0.43±0.21 0.86±0.14 0.63*±0.07
Table 6: Effectiveness F1 scores (± std) for all models. Statistically significant results are marked with *.
Model Twitter Dataset CONAN Dataset
Emo. Aud. Clarity Evid. Rebut. Fair. AVG Emo. Aud. Clarity Evid. Rebut. Fair. AVG
bert_cs 0.31 $\pm$ 0.04 1.00 $\pm$ 0.00 0.52 $\pm$ 0.11 0.91 $\pm$ 0.03 0.78 $\pm$ 0.06 0.54 $\pm$ 0.09 0.68 $\pm$ 0.02 0.61 $\pm$ 0.15 0.99 $\pm$ 0.00 0.98 $\pm$ 0.01 0.97 $\pm$ 0.01 0.97 $\pm$ 0.01 0.99 $\pm$ 0.00 0.92 $\pm$ 0.03
bert_cs_hs 0.30 $\pm$ 0.04 1.00 $\pm$ 0.00 0.49 $\pm$ 0.16 0.91 $\pm$ 0.04 0.77 $\pm$ 0.07 0.48 $\pm$ 0.07 0.66 $\pm$ 0.02 0.69 $\pm$ 0.13 0.99 $\pm$ 0.00 0.97 $\pm$ 0.01 0.97 $\pm$ 0.02 0.88 $\pm$ 0.19 0.97 $\pm$ 0.03 0.91 $\pm$ 0.05
multi-task_d. 0.51 $\pm$ 0.00 1.00 $\pm$ 0.00 0.45 $\pm$ 0.00 0.97 $\pm$ 0.00 0.62 $\pm$ 0.00 0.49 $\pm$ 0.00 0.67 $\pm$ 0.00 0.95 $\pm$ 0.00 1.00 $\pm$ 0.00 0.98 $\pm$ 0.00 0.96 $\pm$ 0.00 0.94 $\pm$ 0.00 0.95 $\pm$ 0.00 0.96 $\pm$ 0.00
multi-task_u. 0.47 $\pm$ 0.00 1.00 $\pm$ 0.00 0.31 $\pm$ 0.00 0.97 $\pm$ 0.00 0.60 $\pm$ 0.00 0.45 $\pm$ 0.00 0.63 $\pm$ 0.00 0.91 $\pm$ 0.00 1.00 $\pm$ 0.00 0.96 $\pm$ 0.00 0.94 $\pm$ 0.00 0.96 $\pm$ 0.00 0.96 $\pm$ 0.00 0.96 $\pm$ 0.00
dependency_m._3e 0.45 $\pm$ 0.00 1.00 $\pm$ 0.00 0.32 $\pm$ 0.00 0.97 $\pm$ 0.00 0.60 $\pm$ 0.00 0.51 $\pm$ 0.00 0.64 $\pm$ 0.00 0.86 $\pm$ 0.00 1.00 $\pm$ 0.00 0.97 $\pm$ 0.00 0.95 $\pm$ 0.00 0.96 $\pm$ 0.00 0.98 $\pm$ 0.00 0.95 $\pm$ 0.00
dependency_m._6e 0.59 $\pm$ 0.00 1.00 $\pm$ 0.00 0.41 $\pm$ 0.00 0.97 $\pm$ 0.00 0.60 $\pm$ 0.00 0.61 $\pm$ 0.00 0.70 $\pm$ 0.00 0.95 $\pm$ 0.00 1.00 $\pm$ 0.00 0.98 $\pm$ 0.00 0.98 $\pm$ 0.00 0.98 $\pm$ 0.00 0.99 $\pm$ 0.00 0.98 $\pm$ 0.00
Table 7: F1 scores per effectiveness dimension on Twitter and CONAN datasets with CS only embeddings. AVG: macro-average. Emo.: Emotional Appeal, Aud.: Audience Adaptation, Clarity: Clarity of Arguments, Evid.: Evidence, Rebut.: Rebuttal, Fair.: Fairness.
Model Twitter Dataset CONAN Dataset
Emo. Aud. Clarity Evid. Rebut. Fair. AVG Emo. Aud. Clarity Evid. Rebut. Fair. AVG
bert_cs 0.31 $\pm$ 0.04 1.00 $\pm$ 0.00 0.52 $\pm$ 0.11 0.91 $\pm$ 0.03 0.78 $\pm$ 0.06 0.54 $\pm$ 0.09 0.68 $\pm$ 0.02 0.61 $\pm$ 0.15 0.99 $\pm$ 0.00 0.98 $\pm$ 0.01 0.97 $\pm$ 0.01 0.97 $\pm$ 0.01 0.99 $\pm$ 0.00 0.92 $\pm$ 0.03
bert_cs_hs 0.30 $\pm$ 0.04 1.00 $\pm$ 0.00 0.49 $\pm$ 0.16 0.91 $\pm$ 0.04 0.77 $\pm$ 0.07 0.48 $\pm$ 0.07 0.66 $\pm$ 0.02 0.69 $\pm$ 0.13 0.99 $\pm$ 0.00 0.97 $\pm$ 0.01 0.97 $\pm$ 0.02 0.88 $\pm$ 0.19 0.97 $\pm$ 0.03 0.91 $\pm$ 0.05
multi-task_d. 0.54 $\pm$ 0.19 1.00 $\pm$ 0.00 0.36 $\pm$ 0.24 0.97 $\pm$ 0.07 0.62 $\pm$ 0.25 0.53 $\pm$ 0.15 0.67 $\pm$ 0.05 0.94 $\pm$ 0.03 0.99 $\pm$ 0.00 0.96 $\pm$ 0.02 0.95 $\pm$ 0.02 0.94 $\pm$ 0.02 0.98 $\pm$ 0.03 0.95 $\pm$ 0.01
multi-task_u. 0.47 $\pm$ 0.19 1.00 $\pm$ 0.00 0.31 $\pm$ 0.12 0.97 $\pm$ 0.07 0.56 $\pm$ 0.19 0.36 $\pm$ 0.14 0.61 $\pm$ 0.07 0.89 $\pm$ 0.02 0.99 $\pm$ 0.00 0.91 $\pm$ 0.03 0.94 $\pm$ 0.02 0.96 $\pm$ 0.02 0.97 $\pm$ 0.01 0.92 $\pm$ 0.01
dep._m._3e 0.38 $\pm$ 0.14 1.00 $\pm$ 0.00 0.37 $\pm$ 0.07 0.97 $\pm$ 0.07 0.61 $\pm$ 0.19 0.48 $\pm$ 0.18 0.63 $\pm$ 0.06 0.84 $\pm$ 0.10 0.99 $\pm$ 0.00 0.96 $\pm$ 0.02 0.93 $\pm$ 0.03 0.94 $\pm$ 0.04 0.96 $\pm$ 0.04 0.94 $\pm$ 0.02
dep._m._6e 0.53 $\pm$ 0.16 1.00 $\pm$ 0.00 0.44 $\pm$ 0.20 0.97 $\pm$ 0.07 0.66 $\pm$ 0.21 0.56 $\pm$ 0.15 0.69 $\pm$ 0.03 0.95 $\pm$ 0.03 0.99 $\pm$ 0.00 0.98 $\pm$ 0.02 0.99 $\pm$ 0.01 0.98 $\pm$ 0.02 0.99 $\pm$ 0.02 0.98 $\pm$ 0.01
Table 8: F1 scores per effectiveness dimension on Twitter and CONAN datasets with CS+HS embeddings. AVG: macro-average. Emo.: Emotional Appeal, Aud.: Audience Adaptation, Clarity: Clarity of Arguments, Evid.: Evidence, Rebut.: Rebuttal, Fair.: Fairness.
Appendix C Statistical Significance Test
The dependency_matrix_6e model consistently achieves the highest average F1 scores across five runs, outperforming all other configurations. A one-way ANOVA test, conducted on the average F1 scores across models, confirms a statistically significant difference in performance (F = 12.92, p < 0.001). To further assess whether dependency_m._6e is significantly better than the rest, we conduct pairwise t-tests between it and each of the other models. Results show that dependency_m._6e significantly outperforms bert_cs (p = 0.000114), bert_cs_hs (p = 0.000010), multi_task_d. (p = 0.003057), multi_task_u. (p = 0.000145), and dependency_m._3e (p = 0.000028) at the 1% significance level. These findings demonstrate that dependency_m._6e provides a statistically significant improvement over other baselines and variants.
Appendix D In-Data Results
Table 7 reports F1 classification results of model configurations, with training and testing on the same dataset for the Twitter Dataset, and for CONAN. The results are the average of the scores obtained from 5 runs with seeds of 42, 0, 1, 2, 3, with CS only embeddings, as the results in the main paper. Table 8 reports results with CS + HS embeddings. In both settings, we confirm the trend presented in the main paper, with the Dependency_matrix_6e achieving the best average F1 score. Furthermore, results show that models trained with only CS embeddings achieve slightly better scores than the ones trained on CS and HS embeddings. This can be due to the fact that adding more context, in this scenario, does not help the model to perform better, but it adds confusion. We also notice that training and testing on expert-written CS (CONAN) seems easier as it achieves higher average F1 scores, compared to training and testing on the Twitter dataset.
Appendix E Error Analysis
In this section, we discuss some examples to better understand the models’ behaviors, in particular on the most difficult class (emotional appeal) and the one achieving almost perfect F1 score (audience adaptation).
Audience Adaptation. It consistently achieves high F1 scores (0.99–1.00) across models, likely due to the similarity in linguistic structure and content between CS and HS. For example, consider the following HS/CS pair from CONAN:
HS: "Islam is not a religion that professes peace and it does not make any good men. It condemns women and encourages delinquency. Lawbreakers are not all of a certain race or religion, but only one race and religion make lawbreakers."
CS: "Today, in many Muslim-majority communities, women receive equal, even higher in some places, treatment as men. For example, many women work as scientists, engineers, lawyers, etc." Both speeches aim to persuade a specific group by matching linguistic complexity, thus ensuring audience adaptation. The CS provides counter-examples to the HS claims, using similarly complex and detailed arguments. This makes it easier for the classifier to recognize this dimension, as both speeches maintain a similar linguistic and argumentative structure. However, this near-perfect performance could reflect data bias, suggesting that audience adaptation might be too easily detected, due to the high class imbalance. Further error analysis should investigate whether models are truly understanding adaptation strategies or simply leveraging surface-level cues.
Emotional Appeal. Emotional appeal is challenging to classify due to its subjective nature and context-dependent expression. For example, consider the following counter-speeches from Twitter and CONAN:
Twitter: "@user this isn’t funny at all. don’t make jokes of the situation when rene herself is affected by it. don’t get her dragged even more and delete this." This CS is direct and uses imperatives to express frustration and urgency, relying on internet culture and social cues like tagging someone and using second-person pronouns ("you").
CONAN: "I thought we’d stopped exporting our convicted criminals last century. Now you advocate exporting people without a trial or a conviction." This example uses sarcasm, rhetorical questioning, and historical references to evoke moral outrage and challenge the opposing argument. The emotional appeal here is more subtle and indirect, contrasting with the explicit frustration in the Twitter example. A model trained solely on Twitter might miss these subtleties found in expert-written CS. To improve emotional appeal classification, domain adaptation or dataset augmentation could help models recognize diverse emotional expressions.